OpenStudy (anonymous):

Solve: −4(x + 1) − 3 = −3(x − 4) −19 19 2 0

OpenStudy (anonymous):

x=-19

OpenStudy (anonymous):

-4x-4-3=-3x+12 Take all the terms with x's to one side: -4x-7=-3x+12 -x=19 x=-19